OpenStudy (anonymous):

solve (8x10^-4)^6 without using a calculator. Answer have to be in scientific notation.

OpenStudy (unklerhaukus):

\[(8\times10^{-4})^6\]\[=8^6\times10^{-4\times6}\]\[=\]

OpenStudy (unklerhaukus):

\[=(2^3)^6\times10^{-4\times6}\] \[=2^{18}\times10^{-4\times6}\]

OpenStudy (radar):

Maybe it was meant to be \[8\times10^{-24}\] but who knows it might be\[8^{6}\times10^{-24}\]

OpenStudy (phi):

using 2^10 = 1024 apprx 1000 or 10^3 we get \[ 256 \cdot 10^{3-24} \approx 2.6 \cdot 10^{-19} \]

OpenStudy (unklerhaukus):

\[=2^{10}\times2^8\times10^{-4\times6}\]\[=1024\times256\times10^{-4\times6}\]\[=(1000\times256+20\times256+4\times256)\times10^{-4\times6}\]\[=(256000+5120+1024)\times10^{-4\times6}\] \[=(2.56000+0.05120+0.01024)\times10^5\times10^{-4\times6}\] \[=(2.56000+0.05120+0.01024)\times10^{-4\times6+5}\] \[=\cdots\]